OpenStudy (anonymous):

As a salesperson at Roaring Waves Beach Supplies, Carissa receives a monthly base pay plus commission on all that she sells. If she sells $500 worth of merchandise in one month, she is paid $385. If she sells $1,000 of merchandise in one month, she is paid $470. Find Carissa's salary when she sells $1,900 worth of merchandise.

OpenStudy (anonymous):

y = 0.17x + 300 so at 1900 she earns $623

OpenStudy (anonymous):

let S= salary per month (constant) and x=% of commission then her pay for each two first month would be: S+500x=385 S+1000x=470 solve it for x - to find % of commission received

OpenStudy (anonymous):

you can solve the system of equation by elimination method = subtract first equation from the second; you'll get: 400x=85 x=0.17 (17%) then put the value of x in any of your original equations (system of two) & find monthly salary: S=385-500*0.17=300 (I used first one) So, her pay for 3d month would be: P=300+1900*.17=623 ($)

OpenStudy (anonymous):

when you are answering this question; what tools may you use? Calculator? then you have to write the equation as shown by inik Graphing Calculator? plot point in calculator and ask calculator for the equation of the line Pencil and paper? as shown by inik Computer? use excel in scatter plot then add trendline(make sure you have x and y values in correct place)